OpenStudy (anonymous):

Is it possible to find the y of x^2 + xy +y^2=3?

OpenStudy (anonymous):

do you mean solve for y=?s if you do you would have to subtract x^2 then divide by x adding like terms geting you to y^3= x^2+3/x then you would just take the cubic root leaving you with y= the cbrt (i don't know how to type one) x^2+3/x hope this helps!- Ginny :)

OpenStudy (anonymous):

Thanks, it helps. I skipped advanced algebra and went right to calc and I have trouble with some of the most basic things.

OpenStudy (anonymous):

wait, after you subtract x^2 wouldn't you get xy + y^2 = 3 - x^2?

OpenStudy (anonymous):

lol you sound like my best friend, she skipped algebra 2 and when to precalc! And I am crazy enough of planning to skip ab calc to get to bc my algebra 2 teacher was aweful i have a link that might help you http://webdesign2.flvs.net/math/precalc.html try this out! :)

OpenStudy (anonymous):

oops yes you caught me! sorry about that long day after the gym :P :)

OpenStudy (anonymous):

 there! :)